# Simulating Relational Event Histories - Why and How

### v2

@@ 5.3 Evaluating Intervention Targeting Strategies
The interventions in our example are carried out by modifying the department-heterophily parameter $\beta_{het}$ only for the selected targets. Specifically, the departmental heterophily parameter during the intervention remains the same as its value pre-intervention ($\beta_{het}^{pre}=-0.58$) for dyads where the sender does not belongs to the target group of actors $\mathcal{Q}$. However, the heterophily parameter is increased to a higher value ($\beta_{het}^{intv}=0.5$) for dyads where the sender belongs to the target group of actors. Thus, the heterophily effect is split into two groups with two different parameter values.
$\centering\beta_{het}(i,j)=\begin{cases}\beta_{het}^{pre}&\ i\notin\mathcal{Q}^{K}\\ \beta_{het}^{intv}&\ i\in\mathcal{Q}^{K}\end{cases}\@add@centering$
@@ 6 Discussion
In this paper, we presented a flexible framework for simulating dynamic networks. Relational event modeling simulations make it possible to study complex interaction dynamics in isolation as well as in combination with other mechanisms that may operate in a dynamic network. Such simulations are effective for a variety of purposes. In this paper, we have explored the reasons why simulating relational event histories is such a valuable tool for social scientists and researchers across disciplines. They can be used for goodness of fit analyses of models fitted on empirical data. They can provide insights into building theories, testing assumptions, evaluating alternative explanations and hypotheses about the operation of interaction mechanisms in dynamic networks. Furthermore, they can be used to evaluate consequences of network interventions,identify intervention points, and inform policy decisions.
@@
The simulation frameworks provide a practical means of simulating dynamic networks that resemble real dynamic networks. This approach while promising, also has certain limitations. First, the task of specifying an appropriate simulation model is not trivial and without enough theory, understanding, or extensive exploration, simulated networks can be prone to mis-specification. Simulations from a mis-specified model may result in unrealistic networks such as networks with unrealistically high (or low) connectivity. However, improving the model specification can help circumvent this problem. Second, the simulations we outlined in our applications only provide control over the local mechanisms of interaction. These local ’rules’ give rise to global network level properties indirectly. However, our simulations do not give direct control over network level properties. For example, if a network of a specific density or degree distribution is desired, then the methods proposed in this paper may not be helpful as the density or degree distribution of a relational event network may only be partially dependent on the effects used to simulate relational event sequences. Other simulation models may be more suited towards these tasks.
@@
Worthwhile directions of future research include relational event simulations methods that take into account heterogeneity in network data, for instance through multilevel models or latent variable models; simulations that incorporate more complex shapes of memory; the possibility to allow network effects to changes gradually or periodically (for example to mimic seasonal behavior); and the development of more advanced goodness of fit indices. The simulation objectives and examples discussed in this article along with the easy to use open source R package [package name redacted for double blind review] can provide the reader with a tool that opens up new avenues of research and shortens the design time for future attempts for simulating time-stamped relational event data.
@@ Acknowledgement
This research was supported by the Netherlands Organization for Scientific Research (NWO) Grant to JM and RL (452-17-006) and an ERC Starting Grant “TIMEISNOW” to JM and RL (758791).

### v4

@@ 5.3 Evaluating Intervention Targeting Strategies
The interventions in our example are carried out by modifying the department-heterophily parameter $\boldsymbol{\beta}(t)_{het}$ only for the selected targets. Specifically, the departmental heterophily parameter during the intervention remains the same as its value pre-intervention ($\boldsymbol{\beta}(t)_{het}^{pre}=-0.58$) for dyads where the sender does not belongs to the target group of actors $\mathcal{Q}$. However, the heterophily parameter is increased to a higher value ($\boldsymbol{\beta}(t)_{het}^{intv}=0.5$) for dyads where the sender belongs to the target group of actors. Thus, the heterophily effect is split into two groups with two different parameter values.
$\centering\boldsymbol{\beta}(t)_{het}(i,j)=\begin{cases}\boldsymbol{\beta}(t)_{het}^{pre}&\ i\notin\mathcal{Q}^{K}\\ \boldsymbol{\beta}(t)_{het}^{intv}&\ i\in\mathcal{Q}^{K}\end{cases}\@add@centering$
@@ 6 Discussion
Relational event simulation techniques are indispensable tools for social network researchers for numerous purposes such as making predictions, building theories, testing model assumptions, evaluating hypotheses on network dynamics, or checking statistical power. To apply these techniques in a relatively easy manner, the paper presented flexible simulation frameworks under dyadic and actor-oriented relational event models which are readily available using the R package remulate. The usefuless of simulation techniques using remulate was shown for different types of problems in social network research. First, the package was used for assessing the goodness of fit of relational event models to study violent interactions between criminal gangs. Simulations indicated that while the model fits well in general, it may underestimate interevent time between escalations in gang conflicts, suggesting a need for model refinement to better capture these dynamics. Second, the simulation framework was used for building, evaluating, and extending social network theories, one of the most important goals of social research. This was demonstrated through the application of the Optimal Distinctiveness Theory, illustrating how individual preferences for inclusion and distinctiveness drive group formation in networks. Third, the simulation framework was used for planning network interventions in real-life, identify intervention points, and inform policy decisions. For example, simulations helped design interventions to increase inter-department communication in an organization, evaluating the persistence and impact of these intervention outcomes over time.
@@
Although the flexible frameworks and software provide a practical means for social network researchers to simulate realistic dynamic networks for different purposes, the application of the methodology remains challenging. First, the task of specifying appropriate simulation models is not trivial and time-consuming. For this reason the use of existing theories (even on a general level), contextual understanding, and extensive exploration is essential to avoid model mis-specification. Simulations from a mis-specified model may result in unrealistic networks such as networks with unrealistically high (or low) connectivity. Network simulation techniques are indispensable to circumvent this problem. Second, the simulations we outlined in our applications only provide control over the local mechanisms of interaction. These local ’rules’ give rise to global network level properties indirectly. However, our simulations do not give direct control over global network level properties. For example, if a network of a specific density or degree distribution is desired, then the methods proposed in this paper may not be helpful as the density or degree distribution of a relational event network only partially depends on the effects used to simulate relational event sequences. Despite these challenges, statistical simulation techniques belong still to the best and most flexible methods of a social network researchers’ toolkit to study social interaction dynamics in complex real-life applications.
@@
Worthwhile directions of future research on temporal social network simulators include relational event models that take into account more complex heterogeneity structures in network data, for instance through crossed effects models or latent variable models, more complex shapes of memory decay in endogenous statistics, and further refinement of time-sensitive goodness of fit indices. The simulation objectives and examples discussed in this article along with the easy to use open source R package remulate can provide the reader with a tool that opens up new avenues of research and shortens the design time for simulating time-stamped relational event data.
@@ Acknowledgement
The first author was supported by a NWO Grant (452-17-006). The second author was partially supported by a NWO Grant (452-17-006) and an ERC Starting Grant (758791). The third author was supported by an ERC Starting Grant (758791).
Conflict of Interest
On behalf of all authors, the corresponding author states that there is no conflict of interest.
Data Availability Statement
Data utilized in Section 3.2, 5.2 and 5.3 are not publicly available. The simulation scripts are available upon request.
Appendix A R Code for simulations using remulate R package
1library(devtools)
2install_github("TilburgNetworkGroup/remulate")
3
4#load the package
5library(remulate)
6
7# define number of actors for the simulations
8N <- 10
9# create a covariates data.frame that contains covariates for each actor in the network
10cov <- data.frame(id=1:N,time=rep(0,N),gender=sample(c(0,1),N,replace=T,prob=c(0.4,0.6)),age=sample(20:30,N,replace=T))
11
12#formula to specify statistics for tie-oriented model
13effects <- ~ remulate::baseline(-3)+
14 remulate::inertia(0.10,scaling="std") +
15 remulate::reciprocity(-0.04)+
16 remulate::same(0.02,variable="gender", attributes = cov)+
17 remulate::send(0.01, variable="age", attributes = cov)+
18 remulate::interact(param=.01, indices = c(2,5)) #interaction of inertia (2) and send age (5)
19
20#generate 100 events using tie oriented model or until time reaches a value of 100
21dat <- remulate::remulateTie(effects = effects, actors = 1:N, events = 100, time = 100)
22
23#formulae to specify statistics for actor-oriented model
24rate_effects <- ~ remulate::baseline(-3) +
25 remulate::outdegreeSender(0.1)+
26 remulate::send(0.01,variable="age",attributes=cov)
27
28choice_effects <- ~remulate::inertia(0.1) +
29 remulate::reciprocity(-0.3)+
30 remulate::same(0.02,variable="gender",attributes = cov)
31
32#generate using actor oriented model
33dat <- remulate::remulateActor(rate_effects, choice_effects, actors = 1:N, events = 100, time = 100)
Appendix B R code of simulations for Section 3.2
1library(remulate)
2library(remstats)
3library(remstimate)
4
5reh <- remify::remify(edgelist = edgelist, model = "tie")
6
7statistics <- ~ remstats::send(variable = "IsPubHouse", attr_actors = actor_cov)+
8 remstats::send(variable = "BonPow", attr_actors = actor_cov)+
9 remstats::send(variable = "size_medium", attr_actors = actor_cov)+
10 remstats::send(variable = "size_large", attr_actors = actor_cov)+
11 remstats::send(variable = "size_verylarge", attr_actors = actor_cov)+
12 remstats::receive(variable = "IsPubHouse", attr_actors = actor_cov)+
13 remstats::receive(variable = "BonPow",attr_actors = actor_cov)+
14 remstats::receive(variable = "size_medium", attr_actors = actor_cov)+
15 remstats::receive(variable = "size_large", attr_actors = actor_cov)+
16 remstats::receive(variable = "size_verylarge", attr_actors = actor_cov)+
17 remstats::tie(attr_dyads = dyad_cov, variable = "distance",scaling="std")+
18 remstats::tie(attr_dyads = dyad_cov, variable = "rivalry")+
19 remstats::inertia(scaling = "prop")+ #FrPSndSnd
20 remstats::rrankReceive() + #RRecSnd
21 remstats::rrankSend() +#RSndSnd
22 remstats::itp(scaling="std")+
23 remstats::otp(scaling="std")+
24 remstats::psABBA()+
25 remstats::psABBY()+
26 remstats::psABAY()
27
28statistics_array <- remstats::tomstats(effects = statistics, reh = reh)
29
30# Estimate the model
31est <- remstimate::remstimate(reh = reh, stats = statistics_array, model = "tie", method = "MLE", ncores = 1)
32
33params <- est\$coefficients
34
35# Simulate from this model
36sim_statistics <- ~ remulate::baseline(param = params["baseline"])+
37 remulate::send(param = params["send_BonPow"], variable = "BonPow", attributes = actor_cov)+
38 remulate::send(param = params["send_size_medium"], variable = "size_medium", attributes = actor_cov)+
39 remulate::send(param = params["send_size_large"], variable = "size_large", attributes = actor_cov)+
40 remulate::send(param = params["send_size_verylarge"], variable = "size_verylarge", attributes = actor_cov)+
41 remulate::receive(param = params["receive_IsPubHouse"], variable = "IsPubHouse", attributes = actor_cov)+
42 remulate::receive(param = params["receive_BonPow"], variable = "BonPow",attributes = actor_cov)+
43 remulate::receive(param = params["receive_size_medium"], variable = "size_medium", attributes = actor_cov)+
44 remulate::receive(param = params["receive_size_large"], variable = "size_large", attributes = actor_cov)+
45 remulate::receive(param = params["receive_size_verylarge"], variable = "size_verylarge", attributes = actor_cov)+
46 remulate::dyad(param = params["tie_distance"], attributes = dyad_cov, variable = "distance",scaling="std")+
47 remulate::dyad(param = params["tie_rivalry"], attributes = dyad_cov, variable = "rivalry")+
48 remulate::inertia(param = params["inertia"], scaling = "prop")+ #FrPSndSnd
49 remulate::rrankReceive(param = params["rrankReceive"] ) + #RRecSnd
50 remulate::rrankSend(param = params["rrankSend"] ) + #RSndSnd
51 remulate::itp(param = params["itp"], scaling="std")+
52 remulate::otp(param = params["otp"], scaling="std")+
53 remulate::psABBA(param = params["psABBA"] )+
54 remulate::psABBY(param = params["psABBY"] )+
55 remulate::psABAY(param = params["psABAY"] )
56
57
58
59# Simulate 100 relational event histories
60net_folder = "./net/"
61runs = 100
62for(r in 1:runs){
63 dat <- remulate::remulateTie(effects = sim_statistics, actors = 1:33, time = 100000, events = nrow(edgelist))
64 save(dat,file=paste0(net_folder,"dat_",r,".rdata"))
65}